# AnTKV: Anchor Token-Aware Sub-Bit Vector Quantization for KV Cache in Large Language Models

### v1

@@ Abstract
Quantization has emerged as an effective and lightweight solution to reduce the memory footprint of the KV cache in Large Language Models (LLMs). Nevertheless, minimizing the performance degradation caused by ultra-low-bit KV cache quantization remains a significant challenge. We observe that quantizing the KV cache of different tokens has varying impacts on the quality of attention outputs. To systematically investigate this phenomenon, we perform forward error propagation analysis on attention and propose the Anchor Score (AnS) that quantifies the sensitivity of each token’s KV cache to quantization-induced error. Our analysis reveals significant disparities in AnS across tokens, suggesting that preserving a small subset with full precision (FP16) of high-AnS tokens can greatly mitigate accuracy loss in aggressive quantization scenarios. Based on this insight, we introduce AnTKV, a novel framework that leverages Anchor Token-aware Vector Quantization to compress the KV cache. Furthermore, to support efficient deployment, we design and develop a triton kernel that is fully compatible with FlashAttention, enabling fast online Anchor Token selection. AnTKV enables LLaMA-3-8B to handle context lengths up to 840K tokens on a single 80GB A100 GPU, while achieving up to 3.5x higher decoding throughput compared to the FP16 baseline. Our experiment results demonstrate that AnTKV matches or outperforms prior works such as KIVI, SKVQ, KVQuant, and CQ under 4-bit settings. More importantly, AnTKV achieves significantly lower perplexity under ultra-low-bit quantization on Mistral-7B, with only 6.32 at 1-bit and 8.87 at 0.375-bit, compared to the FP16 baseline of 4.73.
@@ 1 Introduction
LLMs have attracted much attention due to their strong capability in natural language processing and other artificial intelligence tasks openai2023gpt4 ; guo2025deepseek ; team2023gemini ; tang2025the ; dong2025can . With the rapid development in recent years, LLM can manage increasingly larger context lengths ranging from hundreds of thousands to millions of tokens, and its broadening scope allows LLM to tackle more complex tasks zhao2023survey ; wang2025all ; das2025security ; wang2025agenttaxo . However, these models, which are typically built on decoder-based transformer architectures, generate tokens recursively during inference, thus leading to a large accumulation of the KV cache as the context length zhu2025oraclekv ; bai2024longbench ; liu2025comprehensive and batch size increases pope2022efficientlyscalingtransformerinference . The large memory footprint of the KV cache presents a significant challenge. For long-context inference, such as LLaMA-3’s 128K context length, the KV cache can consume several gigabytes of memory, nearly matching the model’s weights. This challenge is even greater for models like Gemini, which handle contexts up to a million tokens. Moreover, LLM-RAG systems yao2025cacheblend pre-generate and store KV caches. However, the excessive size of these caches presents a significant challenge for storage.
@@
Figure 1: Self-attention output error when quantizing the KV cache of Mistral-7B to 1-bit.
@@
Figure 2: The perplexity of Mistral-7B on the Wikitext2 under different bit-widths.
@@
To address the growing KV cache, various optimization strategies have been proposed liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; liu2024retrievalattention ; zhang2024cam ; liu2024kivi ; dong2025can , with quantization emerging as a prominent solution. Quantization approaches fall into two main categories: scalar quantization (SQ) and vector quantization (VQ). SQ compresses the KV cache by mapping floating-point values to fixed low-bit representations liu2024kivi ; duanmu2024skvq ; hooper2024kvquant , but the minimum 1-bit requirement limits its compression to 1/16 on FP16. In contrast, VQ maps high-dimensional vectors to a finite codebook lingle2024transformervq ; zhang2024kv , capturing intra-vector correlations. This enables higher compression rates and better preserves representational fidelity, especially under ultra-low-bit regimes.
During the LLM inference, different tokens have different effects on the accuracy of LLMs, so it is unreasonable to treat all tokens equally in VQ methods. To illustrate this, we quantize the KV cache of Mistral-7B mistral to 1-bit using VQ, and Figure 1 shows the impact of KV quantization on the self-attention output error within layer 31. It reveals that even in the absence of tokens with exceptionally large quantization errors, a few tokens (Anchor Tokens) can still introduce substantial deviations in the output of the self-attention operator. Therefore, measuring the importance of tokens during inference can further explore the performance potential of VQ methods for compressing the KV cache.
To this end, we introduce Anchor Score (AnS), a metric derived from forward error analysis, to quantify each token’s impact on the output. Leveraging AnS, we propose AnTKV, a novel framework that employs anchor token-aware VQ for KV cache compression. AnTKV mitigates quantization-induced degradation through a two-stage design: offline token-aware centroids learning and online anchor token selection. In the offline stage, the framework performs the token-aware weighted k-means clustering to generate centroids. During the online stage, it computes the AnS for each request to identify anchor tokens that significantly contribute to output error and selectively retains their FP16. Moreover, to ensure efficiency, we expose key intermediate variables from FlashAttention dao2023flashattention2 and integrate a lightweight GPU kernel for AnS computation during inference, enabling efficient runtime anchor token selection.
@@
To validate AnTKV’s effectiveness, we conduct extensive experiments across various quantization bit-widths on a range of LLMs, including LLaMA-2/3 llama2 ; llama3 and Mistral-7B mistral family. As depicted in Figure 2, AnTKV consistently achieves better performance across a wide spectrum of bit-widths, from 4-bit down to 0.375-bit. On Wikitext2 with Mistral-7B, AnTKV achieves a perplexity of 6.32 at 1-bit, a reduction of 0.93 compared to CQ (7.25) zhang2024kv , and substantial 9.04 over KVQuant-1% (15.36) hooper2024kvquant . Even at an aggressive 0.375-bit, AnTKV further lowers perplexity to 8.87, surpassing CQ (13.20) by 4.33. Enabled by the efficient implementation of AnS, AnTKV scales to extremely long contexts. On LLaMA-3-8B, it supports up to 840K context length under 0.375-bit on a single A100-80GB GPU. During decoding, AnTKV increases the maximum batch size by 3.3x and improves throughput by 3.4x at 1K context length compared to FP16. More experiments that demonstrate the performance advantages of the proposed AnTKV are in Section 4.
@@
  • We conduct an error analysis of the self-attention, and introduce AnS to quantify the importance of tokens. With AnS, we can identify anchor tokens that significantly impact the accuracy of LLMs. Based on AnS, we present a novel token-aware VQ method, namely AnTKV, to compress the KV cache in LLM inference.
  • To ensure efficiency, we design a custom GPU kernel for computing AnS in the online stage that is compatible with FlashAttention. With the high-performance implementation of AnS, AnTKV can support up to 840K context length on a single GPU.
  • We deploy the proposed AnTKV in LLaMA-2/3 and Mistral-7B family, numerical results show that AnTKV has a significant performance improvement under various quantization bit-widths.
@@ 2.1 Transformer and Attention
The transformer block has become the fundamental architecture of LLMs; it consists of a fully connected feed-forward layer and a self-attention layer. The self-attention layer enables the model to capture connections among tokens within the context and serves as the critical component of the transformer block. Specifically, it maps a query ($\bm{Q}$) and a set of KV pairs ($\bm{K}$ and $\bm{V}$) to an output, i.e.,
@@ 2.2 Memory Constraints in LLM Inference
During the process of LLM inference, the key and value generated at prefill and each decoding step are stored to avoid redundant computation. The KV cache is repeatedly accessed in subsequent steps to compute attention over the full context seen so far, significantly accelerating autoregressive decoding. In a multi-turn conversation, the KV cache from previous turns can also be reused during the prefill stage to further reduce latency. The state-of-the-art open-source and closed-source LLMs, like LLaMA-3.1-8B llama3 and Gemini 1.5 Pro team2023gemini , now handle much longer context lengths, up to 128K and 2 million tokens, respectively. This increase in context length liu2025comprehensive ; wang2025agenttaxo causes the GPU memory used for the KV cache in LLM inference to even exceed the models’ size. The rapid growth of KV cache greatly limits the deployment of LLMs with long context. Quantizing the KV cache directly optimizes this by maintaining key-value information and cutting memory costs substantially, thus enhancing the efficiency of LLM inference.
@@ 2.3 Related Works
KV cache quantization A variety of KV cache quantization methods have been developed to address the memory bottleneck in long-context LLMs liu2024kivi ; duanmu2024skvq ; hooper2024kvquant ; zhang2024kv ; zhu2024sampleattentionnearlosslessaccelerationlong . KIVI liu2024kivi reduces quantization error via per-channel key quantization combined with a sliding window strategy to emphasize locally relevant tokens. SKVQ duanmu2024skvq further explores this direction by introducing channel reordering and clipping to improve low-bit quantization. In contrast, KVQuant hooper2024kvquant adopts a different design by quantizing pre-RoPE keys using non-uniform quantization and selectively preserving a small fraction of full-precision KV entries to mitigate accuracy loss. CQ zhang2024kv adopts a VQ-based approach, aiming to exploit cross-channel correlations to further compress the KV cache.
KV cache compression Beyond quantization, the field of LLMs is actively exporing several advanced methods for KV cache compression. Sparse attention are designed to achieve compression by selectively computing attention over a reduced set of tokens xiaoefficient ; chen2024prefixquant ; zhu2025oraclekv ; liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; li2024snapkvllmknowslooking , However, a significant limitation lies in the inherent complexity of designing and optimizing their sparsity patterns. This often necessitates task-specific or data-driven heuristics, which can severely constrain their efficient generalization across diverse model architectures and datasets. Token merge, as a variant of sparse attention, merges the KV cache of similar tokens during inference to reduce memory usage zhang2024cam ; wang2024modeltellsmergeadaptive . Retrieval-based methods liu2024retrievalattention ; chenmagicpig ; zhang2024pqcache compress KV caches by retrieving relevant historical context for each query and are effective in tasks like RAG or document retrieval. However, they often lose substantial KV information, making them less suitable for general-purpose LLM inference.
@@
Model Compression Many model compression techniques closely align with KV cache compression in both goals and underlying methodologies. GPTQ frantar2023gptqaccurateposttrainingquantization utilizes calibration set to reduce quantization induced degradation, while SmoothQuant xiao2024smoothquantaccurateefficientposttraining and AWQ lin2024awqactivationawareweightquantization minimize output error from the perspective of error propagation analysis. VQ-based methods such as VPTQ liu2024vptqextremelowbitvector and QUIP# tseng2024quipbetterllmquantization further enhance compression fidelity through residual VQ and Hadamard transform. Additionally, system-level frameworks like Atom zhao2024atomlowbitquantizationefficient and QServe lin2025qservew4a8kv4quantizationcodesign jointly optimize model and KV cache quantization, achieving end-to-end acceleration under low-bit quantization. Other works, such as Pruner-Zero dong2024prunerzeroevolvingsymbolicpruning and Parzc dong2024parzcparametriczerocostproxies , explore how to sparsify model weights while preserving model performance. Systems like FlashLLM xia2023flashllmenablingcosteffectivehighlyefficient and Spinfer fan2025spinfer leverage the model sparsity to accelerate inference.
@@ 3 Methodology
Previous works on quantifying the KV cache typically do not distinguish the importance of individual tokens, but assume that all tokens contribute equally to attention accuracy, or they adopt a simpler strategy of only preserving designated sink tokens. However, our 1-bit VQ experiment on Mistral-7B contradicts this, unveiling a previously unaddressed aspect in KV cache quantization: a small subset of tokens provokes disproportionately large perturbations in the self-attention outputs. This divergence reveals a new opportunity for KV cache quantization. Instead of uniform application, the precision of quantization should correspond to each token’s individual importance. Specifically, if we identify and preserve a subset of high-impact tokens in FP16, we can unlock a far better memory–accuracy trade-off.
Inspired by this, we present a novel token-aware VQ method for KV cache compression, namely, AnTKV. In AnTKV, we adopt a two-stage design: offline token-aware centroids learning (see Figure 3(a)) and online anchor token selection (see Figure 3(b)), the key of which lies in how to measure the importance of tokens. In the offline stage, we perform a weighted k-means clustering to produce centroids in VQ, and the gradient corresponding to each token is used as weights. In the online stage, to reduce the output error caused by quantization, we select tokens with both higher $\text{Err}_{\text{quant}}$ and AnS, i.e. anchor tokens, and keep them at FP16. Additionally, since applying RoPE disrupts the consistency of channel-wise magnitudes in $\bm{K}$ (see Appendix A for an illustrative example), which exhibits large inter-cluster distances and small intra-cluster variances, leading to large quantization error, the pre-RoPE strategy with the same as hooper2024kvquant is used in AnTKV.
Figure 3: The overview of AnTKV. (a) In the offline stage, we perform token-aware centroid learning to obtain the centroids. (b) In the online stage, we apply VQ using centroids and compute the AnS. Then identify the top-$k$ anchor tokens with the highest product of VQ error and AnS, and retain them in full-precision to preserve model performance.
@@ 3.1 Offline Token-Aware Centroids Learning
The offline stage aims to learn centroids by aggregating KV cache from general calibrations. Specifically, each row (per token) of $\bm{K}$ and $\bm{V}$ is divided into several sub-vectors, then using k-means to cluster them, and only the centroids are stored. To integrate the importance of tokens, we introduce the weighted k-means instead of classical k-means for the offline centroids learning.
From the perspective of accuracy, we hope that the errors caused by quantizing $\bm{K}$ and $\bm{V}$ will not affect the accuracy of LLMs as much as possible. Mathematically, it can be formulated as
$\begin{split}\min\limits_{\bm{C}^{k}\in\mathbb{R}^{c^{k}\times d}}|L(\bm{K},\bm{V})-L(\texttt{vq}(\bm{K}),\texttt{vq}(\bm{V}))|,\\ \end{split}$ (1)
@@
where $\frac{\partial L}{\partial\bm{K}_{j,:}}$ and $\frac{\partial L}{\partial\bm{V}_{j,:}}$ are the gradient of the loss function $L(\cdot)$ with respect to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$, i.e., the $j$th row of $\bm{K}$ and $\bm{V}$ that correspond to the $j$th token. From (1) and (2), the weighted k-means clustering with gradients as weights is essentially to find a clustering strategy that minimizes the error caused by KV quantization. This also explains the effectiveness of the weighted k-means clustering used in this stage.

### v2

@@ Abstract
Quantization has emerged as an effective and lightweight solution to reduce the memory footprint of the KV cache in Large Language Models. Nevertheless, minimizing the accuracy degradation caused by ultra-low-bit KV cache quantization remains a significant challenge. While scalar quantization is constrained by 1-bit bound, vector quantization exploits intra-vector correlations and enables sub-bit regimes, making it more suitable for ultra-low-bit quantization. To further mitigate quantization-induced degradation, we reveal that the degradation is highly uneven across tokens in attention quality. To investigate this unevenness, we introduce anchor score to measure each token’s sensitivity to quantization. Our analysis and experiments show that preserving a small subset (1%) of tokens with the highest Anchor Score significantly mitigates accuracy loss under aggressive quantization.
We propose AnTKV, a dual-stage framework that leverages anchor token-aware vector quantization to compress the KV cache. It combines offline token-aware centroids learning and online anchor token selection to balance compression and accuracy. To enable efficient deployment, we design an online anchor token selection kernel compatible with FlashAttention. It allows LLaMA3-8B to scale to 840K tokens on a single 80GB A100, while delivering up to $3.5\times$ higher decoding throughput over the FP16 baseline. Experiments demonstrate that AnTKVmatches or surpasses prior methods at 4-bit, and significantly reduce perplexity under ultra-low-bit quantization, achieving 6.32 at 1-bit on Mistral-7B, compared to 7.25 for CQ and 15.36 for KVQuant.
@@ 1 Introduction
Large Language Models (LLMs) have gained wide attention owing to their remarkable capabilities in diverse applications openai2023gpt4 ; guo2025deepseek ; team2023gemini ; tang2025the ; dong2025can . With rapid recent advances, LLMs currently handle context lengths from hundreds of thousands to millions of tokens, enabling them to tackle increasingly complex tasks zhao2023survey ; wang2025all ; das2025security ; wang2025agenttaxo . Most LLMs adopt decoder-based transformer architectures, where tokens are generated autoregressively and the KV cache grows rapidly with context length zhu2025oraclekv ; bai2024longbench ; liu2025comprehensive and batch size increases pope2022efficientlyscalingtransformerinference . The large amount of memory footprint of the KV cache poses a significant challenge. For example, with LLaMA-3 at 128K tokens it already approaches the model size, and for million-token models like Gemini it becomes prohibitive. Beyond inference, LLM-RAG systems yao2025cacheblend pre-generate and store massive KV caches, creating additional storage challenges.
@@
Figure 1: The $L_{1}$ norm error of attention output when quantizing the $i$th token’s KV cache in Mistral-7B to 1-bit.
@@
Figure 2: The Perplexity of Mistral-7B on the WikiText-2 across different quantization bit-widths.
@@
To address the substantial size of KV caches, numerous techniques have been explored liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; liu2024retrievalattention ; zhang2024cam ; liu2024kivi ; dong2025can , with quantization proving to be especially effective. Quantization methods are generally classified into scalar quantization (SQ) and vector quantization (VQ). SQ compresses the KV cache by mapping floating-point values to fixed low-bit representations liu2024kivi ; duanmu2024skvq ; hooper2024kvquant , but the 1-bit lower bound constrains its maximum compression ratio to 1/16 of FP16. In contrast, VQ compresses high-dimensional vectors by mapping them to a finite codebook lingle2024transformervq ; zhang2024kv , which captures intra-vector correlations, achieves superior compression ratios and fidelity under ultra-low-bit quantization.
We observe that tokens contribute unequally to model accuracy during inference. To illustrate this, we quantize the KV cache of Mistral-7B mistral to 1-bit using VQ. Figure 1 depicts the attention output $L_{1}$ norm error in the 31st layer when quantizing the KV cache across different token indices, and similar error distributions are observed in other layers. As evident from the figure, although per-token quantization errors are similar, the resulting error in the attention output vary widely across tokens and a small subset exhibits error that are tens of times larger when quantized (anchor tokens). Quantifying token importance is essential for unlocking the full potential of ultra-low-bit KV cache quantization.
To address this, we conduct a forward error analysis forwarderroranalysis on attention-based models with respect to KV, and then propose AnTKV, a dual-stage framework for KV cache quantization. In the offline stage, AnTKV performs token-aware weighted k-means clustering to generate centroids, where the weights are derived from error propagation factors obtained through forward error analysis. Tokens that cause larger increases in output error are assigned larger weights during clustering. However, the propagation factors entail substantial gradient cost. To overcome this limitation, we propose Anchor Score (AnS) derived from the forward error analysis of the attention, which quantifies the output sensitivity to quantizing each token’s KV cache. At inference time, AnS is computed for each prompt to identify this small token subset.  AnTKV handles this subset of tokens in a simple yet effective manner by preserving them in full precision, thereby reducing accuracy loss. Moreover, we design and implement a lightweight GPU kernel for AnS computation. More specifically, we extend FlashAttention dao2023flashattention2 to store low memory overhead softmax intermediate results, thereby enabling efficient online anchor token selection.
@@
We conduct extensive experiments to evaluate the effectiveness of AnTKV across various quantization bit-widths on a range of LLMs, including the LLaMA-2/3 llama2 ; llama3 and Mistral-7B mistral families. As shown in Figure 2AnTKV consistently outperforms existing approaches across bit-widths from 4-bit down to 0.375-bit. On WikiText-2 wikitext2 AnTKV achieves a perplexity of 6.32 at 1-bit, reducing error by 0.93 compared to CQ (7.25) zhang2024kv , and by a substantial 9.04 compared to KVQuant-1% (15.36) hooper2024kvquant . Even under the aggressive 0.375-bit setting, AnTKV attains a perplexity of 8.87, surpassing CQ (13.20) by 4.33. Thanks to the efficient GPU implementation of AnS, AnTKV also scales to extremely long contexts: on LLaMA3-8B, it supports up to 840K tokens under 0.375-bit quantization on a single A100-80GB GPU. Moreover, during decoding, AnTKV increases the maximum batch size by $3.3\times$ and improves throughput by $3.4\times$ at a 1K context length compared to full precision.
@@
  • To the best of our knowledge, this work is the first to investigate the feasibility of quantization the KV cache to sub-bit while preserving model accuracy.
  • We identify that different tokens contribute unequally to model accuracy under quantization and highlight the existence of anchor tokens that dominate output error.
  • We propose AnTKV, which performs token-aware weighted clustering offline and leverages AnS online to efficiently identify anchor tokens, preserving them in full precision to mitigate accuracy loss, and we evaluate it on the LLaMA2/3 and Mistral families, where it consistently improves performance across a wide range of quantization bit-widths.
  • We implement custom GPU kernels for the online stage, enabling AnTKV to scale to 840K tokens on a single GPU and deliver significant throughput gains during decoding.
@@ 2.1 Transformer and Attention
The transformer block has become the fundamental architecture of LLMs; it consists of a fully connected feed-forward network and an attention. The attention enables the model to capture connections among tokens within the context. Specifically, it maps a query ($\bm{Q}$) and a set of KV pairs ($\bm{K}$ and $\bm{V}$) to an output, i.e.,
@@ 2.2 Memory Constraints in LLM Inference
During the process of LLM inference, the key and value generated at prefill and each decoding step are stored to avoid redundant computation. The KV cache is repeatedly accessed in subsequent steps to compute attention over the full context seen so far, significantly accelerating autoregressive decoding. In a multi-turn conversation, the KV cache from previous turns can also be reused during the prefill stage to further reduce latency. Nowadays LLMs, like LLaMA3.1-8B llama3 and Gemini 1.5 Pro team2023gemini , now handle much longer context lengths, up to 128K and 2 million tokens, respectively. This increase in context length liu2025comprehensive ; wang2025agenttaxo causes the GPU memory consumed by KV cache to even exceed the model. The rapid growth of KV cache greatly limits the deployment of LLMs with long context. Quantization is an effective approach to compressing the KV cache, substantially reducing memory footprint while preserving essential information.
@@ 2.3 Related Works
KV cache quantization A variety of KV cache quantization methods have been proposed to address the memory bottleneck in long-context LLMs liu2024kivi ; duanmu2024skvq ; hooper2024kvquant ; zhang2024kv ; zhu2024sampleattentionnearlosslessaccelerationlong . KIVI liu2024kivi mitigates quantization error by applying per-channel key quantization and employing a sliding window to emphasize locally relevant tokens. SKVQ duanmu2024skvq further explores this direction by introducing channel reordering and clipping. To further reduce accuracy loss, KVQuant hooper2024kvquant introduces pre-RoPE key quantization, non-uniform format and element-wise outlier. CQ zhang2024kv adopts a VQ-based approach, aiming to exploit cross-channel correlations to further compress the KV cache.
KV cache compression Beyond quantization, the field of LLMs is actively exploring advanced methods for KV cache compression. Sparse attention aims to reduce memory footprint by selectively handling the KV cache in a token-wise manner xiaoefficient ; chen2024prefixquant ; zhu2025oraclekv ; liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; li2024snapkvllmknowslooking . However, it discards the KV cache of a subset of tokens, even though the corresponding tokens may be required in subsequent decoding. Token Merging reduces memory usage by consolidating the KV caches of similar tokens during inference, achieving an effect related to sparse attention but through merging rather than dropping tokens zhang2024cam ; wang2024modeltellsmergeadaptive . Retrieval-based methods liu2024retrievalattention ; chenmagicpig ; zhang2024pqcache offload and index KV caches, retrieving a subset of relevant entries for each query, but introduce additional communication overhead.
@@
Model Compression Numerous model compression techniques share common objectives and methodological foundations with KV cache compression. GPTQ frantar2023gptqaccurateposttrainingquantization utilizes calibration set to reduce quantization induced degradation, while SmoothQuant xiao2024smoothquantaccurateefficientposttraining and AWQ lin2024awqactivationawareweightquantization minimize output error from the perspective of error propagation analysis. VQ-based methods such as QUIP# tseng2024quipbetterllmquantization further enhance compression fidelity through Hadamard transform. Pruner-Zero dong2024prunerzeroevolvingsymbolicpruning and Parzc dong2024parzcparametriczerocostproxies , explore how to sparsify model weights while preserving model performance. System-level works like Atom zhao2024atomlowbitquantizationefficient and QServe lin2025qservew4a8kv4quantizationcodesign Recent efforts jointly quantize model, KV cache and activatioin, enabling inference under low-bit and leveraging low-precision Tensor Cores to improve system performance, while approaches such as FlashLLM xia2023flashllmenablingcosteffectivehighlyefficient and Spinfer fan2025spinfer accelerate inference by leveraging model sparsity.
@@ 3 Methodology
VQ is employed as it enables sub-bit compression, which SQ cannot achieve, and in the ultra-low-bit regime ($\leq 2$ bit) it leverages intra-vector correlations to attain significantly better quantization quality compared to SQ.
As shown in the figure 1, we observe that although per-token KV cache quantization errors appear similar, the resulting attention output errors differ dramatically. Moreover, the error distribution is steep, with large errors occurring unpredictably across tokens. Previous works on KV cache quantization typically treat all tokens equally or are biased toward attention sinks, as inspired by AttentionSink xiaoefficient , which we argue is suboptimal. Our observation reveals a new opportunity for KV cache quantization, where tokens with greater influence on attention output error should be prioritized. Building on this observation, we propose AnTKV, a token-aware VQ framework for KV cache quantization. It adopdts a dual-stage design consisting of offline token-aware centroid learning (Figure 3(a)) and online anchor token selection (Figure 3(b)). The central challenge lies in effectively measuring token importance. During centroids leanring, forward error analysis provides per-token error propagation factors, which serve as weights in a token-aware $k$-means clustering to generate VQ centroids, thereby emphasizing tokens with significant influence on attention-output error. For anchor token selection, directly using error propagation factors incurs expensive gradient computation. To avoid this cost, we introduce AnS, a lightweight metric derived from attention forward error analysis that enables efficient prompt-aware identification of anchor tokens.
Figure 3: Overview of AnTKV. In the stage (a), token-aware centroids are learned from calibration data through weighted clustering, where the weights are error-propagation factors obtained by forward error analysis. In the stage (b), the KV cache is quantized with centroids, and AnS is computed to identify anchor tokens, which are preserved in full precision to mitigate accuracy loss.
@@ 3.1 Offline Token-Aware Centroids Learning
In this stage, centroids are learned offline by aggregating KV cache from calibration data. Specifically, each row (per token) of $\bm{K}$ and $\bm{V}$ is divided into several sub-vectors, then using $k$-means to cluster them, and only the centroids are stored. Specifically, each token row of $\bm{K}$ and $\bm{V}$ is partitioned into sub-vectors and clustered with $k$-mean. The centroids are retained for online KV cache quantization.
From an accuracy perspective, replacing $\bm{K}$ and $\bm{V}$ with their corresponding centroids should introduce minimal impact on model output. Mathematically, it can be formulated as
[…]
@@
where $\frac{\partial L}{\partial\bm{K}_{j,:}}$ and $\frac{\partial L}{\partial\bm{V}_{j,:}}$ are the gradient of the loss function $L(\cdot)$ with respect to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$, i.e., the $j$th row of $\bm{K}$ and $\bm{V}$ that correspond to the $j$th token. From equation (1) and (2), the weighted k-means clustering with gradients as weights is essentially to find a clustering strategy that minimizes the error caused by KV quantization.